Solve the initial-value problem. X ( )x 2 4 -16 X, X(0) = (1)please show work
Answer #1

We need 2. 6 To find another eigen vector to generalized Eigen bechor (A-Z2x = Y > ㅉ 거 -AM+4Mg=2 매 월 쿠매 제이 2 멧 / 짓기 사요 2리지 적so X4 - 6 + 13/2te helt te ut + 42 e ut 740) = 126 teht_e4 te This satisfies Ivy o to 4t +6096 final solution is X (t) - [ (2